OpenStudy (anonymous):

Ira bought a tennis racquet that cost $112. The sales tax rate is 9 percent. What is the total amount that she paid? $121 $122.08 $122.80

OpenStudy (johnweldon1993):

So she paid the 112 PLUS whatever 9 percent of that 112 is.. \[\large Total = 112 + (0.09 \times 112) = ?\]

OpenStudy (anonymous):

im tryna put it in the calculator but it doesnt make any sense

OpenStudy (hysusonic):

do you know how to find percentage?

OpenStudy (anonymous):

no not at all thats what has me confused

OpenStudy (hysusonic):

okay so to find the percentage of something, you need to divide the percentage by the number they give you so 9% translates into 0.09 in decimal form

OpenStudy (er.mohd.amir):

first find tax =(9*112)/100 then total cost =112+tax

OpenStudy (hysusonic):

\[0.09/112\] which gives you some number then you multiply that result by 100 and then you add the final result to 112

OpenStudy (anonymous):

but if it comes back to 112 as a result what would i divide or times it by the ???

OpenStudy (anonymous):

because i still doing what your saying but its not giving me one of the answers on the top that i am looking for

OpenStudy (hysusonic):

what?.... what did you get as your final result?

OpenStudy (anonymous):

0.000803571428571

OpenStudy (hysusonic):

wow okay umm you must've done some thing wrong step 1 convert 9% into a decimal which would be 0.09 step 2 divide 0.09 with 112 to get 0.075

OpenStudy (hysusonic):

are you with me so far?

OpenStudy (anonymous):

an i did divide an im getting the same answer

OpenStudy (hysusonic):

why would you divide after this? you're done dividing after step 2

OpenStudy (johnweldon1993):

Where are you getting that 0.075 from? Step 1) Find the tax...9 percent = 0.09....multiply that 0.09 * 112 = some number That is your tax Step 2) Add the tax to the original amount to find the total cost

OpenStudy (anonymous):

thank you i got the answer B